Is it canbe serve or cannot serve, Explain.P(X) = 3 + X for x = 1, 2, 3, 4


3 – X

1
Expert's answer
2022-03-01T08:16:12-0500

"P(X)=\dfrac{3+X}{3-X}" , for X = 1, 2, 3, 4

To check whether P(X) serves a probability distribution.

Solution:

X=1

"P(X)=\dfrac{3+1}{3-1}=\dfrac{4}{2}=2>1"

which is not possible for a probability distribution.

Thus, it does not serve as a probability distribution.